OpenStudy (anonymous):

hat is the radius of a sphere with a surface area of 100πcm^2?

OpenStudy (anonymous):

 find r then by canceling out pi from both sides.....it will be 5cm

OpenStudy (anonymous):

@vidushi, please use the text option if you are going to do a drawing.

OpenStudy (displayerror):

Surface area of a sphere is given by \[4 \pi r^2\] Set that equal to the surface area (which you're given) \[100 \pi = 4 \pi r^2\] \[25 = r^2\] \[r = \pm 5\] Radius can only be positive, so reject the negative 5 value \[r=5\]

OpenStudy (anonymous):

ok rrrrrr

OpenStudy (goldphenoix):

(Sphere) S.A = \[100πcm^2\] To find the surface area of any sphere, you must use the formula: \[4\Pi r^2\]. Since you have the surface area, you must go backward. \[4\pi r^2 = 100\pi cm^2\] Divide 4: \[\pi r^2 = 25 \pi cm^2\] Square root it: \[\pi = 5 \pi cm^2\] Divide pi. R= \[R= 5 CM\]

OpenStudy (anonymous):

thanks everyone

OpenStudy (goldphenoix):

No problem. :)